OpenStudy (anonymous):

Find the number halfway between 1/2 and 5/6

OpenStudy (anonymous):

A. 27/40 B. 2/3 C. 13/20 D. 7/10

OpenStudy (anonymous):

i believe to get the midpoint - which is the same thing? :) you add them and then divide by 2 \[\frac{\frac{ 1 }{ 2 }+\frac{ 5 }{ 6 }}{2}\]

OpenStudy (anonymous):

to add them we should find a common denominator. /2 and /6 definately have /6 as a common (3x /2 = /6

OpenStudy (anonymous):

Get Denominator Joint again !

OpenStudy (anonymous):

so I extend the fraction with 2 in the denominator by 3:\[\frac{\frac{ 1 \times 3}{ 2 \times 3}+\frac{ 5 }{ 6 }}{2}\]and then add them together \[\frac{\frac{ 3}{ 6}+\frac{ 5 }{ 6 }}{2} = \frac{\frac{ 8 }{ 6 }}{2} \]

OpenStudy (anonymous):

have : 5/6 and 3/6 we have 4/6 between them ! 4/6= ?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

Its 2/3?

OpenStudy (anonymous):

Yes !!!

hero (hero):

Just multiply 1/2 by 3/3 to get 3/6 Now you have 3/6 and 5/6 and easy to see that 4/6 is between

hero (hero):

@TakeA.I.M.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

  ^^ For People who are not Fan of Decimals...